One positive number is 5 more than a second​ number, and their product is 150
. Find the two numbers.

Answer by Boreal(15235)

x, and x-5 are the numbers
x*(x-5)=150
x^2-5x-150=0
(x-15)(x+10)=0
x=15 only positive root.
Numbers are 10 and 15.